find all real solutions of the following equations by using the quadratic formula.
1. xSQUARED-6x+8=0
2. xSQUARED-2x=15
3. xSQUARED-8x=-16
4. xSQUARED+10x+25=0
thank you for answering.

Answer by mangeshpjoshi(15)

x^2-6x+8=0
x^2-4x-2x+8=0 try to distribute 6 into such 2 parts so that sum =-6 and multiplication = +8
x(x-4-2(x-4)=0
(x-4)(x-2)=0
x=4 or x=2 is the answer to first question
2)x^2-2x=15
x^2-5x+3x-15=0
x(x-5)+3(x-5)=0
(x-5)(x+3=0
x=5 or x=-3
solution to this
3)x^2-8x+16=0
x^2-4x-4x+16=0
x(x-4)-4(x-4)=0
(x-4)(x-4)=0
x=4 is uniform answer
4)x^2+10x+25=0
x^2+5x+5x+25=0
x(x+5)+5(x+5)=0
(x+5)(x+5)=0
x=-5 uniform soultion